# Customizing Speech Recognition Model with Large Language Model Feedback

### v1

@@ I Introduction
In contrast, large language models (LLMs) such as GPT4[5], Deepseek[6], Phi[7] trained on internet-scale corpora, demonstrate remarkable generalization capabilities across a wide range of domains. These models have demonstrated strong performance on tasks involving commonsense reasoning, in-context learning, and even zero-shot adaptation. Such capabilities suggest that LLMs could serve as valuable resources for enhancing the performance of ASR models, particularly in low-resource or domain-mismatched settings. Recent studies have explored integrating LLMs with ASR systems through approaches such as re-ranking hypotheses[8, 9, 10, 11] or error correction during post-processing[12, 13, 14].
@@
Our experiments demonstrate that our RL framework can substantially improve transcription quality in domain-mismatched scenarios, particularly for named entities and context-sensitive phrases. Compared to standard self-training domain adaptation methods, our approach achieves up to 21% relative improvement on entity word error rate (EWER), showcasing the effectiveness of LLMs feedback in guiding ASR model adaptation.

### v2

@@ I Introduction
In contrast, text large language models (LLMs) such as GPT4[5], Deepseek[6], Phi[7] trained on internet-scale corpora, demonstrate remarkable generalization capabilities across a wide range of domains. These models have demonstrated strong performance on tasks involving commonsense reasoning, in-context learning, and even zero-shot adaptation. Such capabilities suggest that LLMs could serve as valuable resources for enhancing the performance of ASR models, particularly in low-resource or domain-mismatched settings. Recent studies have explored integrating LLMs with ASR systems through approaches such as re-ranking hypotheses[8, 9, 10, 11] or error correction during post-processing[12, 13, 14].
@@
Our experiments demonstrate that our RL framework can substantially improve transcription quality in domain-mismatched scenarios, particularly for named entities and context-sensitive phrases. Compared to standard self-training domain adaptation methods, our approach achieves up to 21% relative improvement on entity word error rate (EWER), showcasing the effectiveness of LLMs feedback in guiding ASR model adaptation. Our contributions can be summarized as follows:
  • We proposed novel reward function for RL in ASR tasks. Unlike prior approaches that rely on a separately trained reward model or human feedback, our method introduces a straightforward reward computation. Furthermore, our reward signal is dense and avoids common issues such as reward hacking, making it both effective and robust.
  • We applied the state-of-the-art RL algorithms to achieve contextual customization for ASR in the unsupervised domain adaptation setting.
  • By leveraging text LLMs, our approach yields up to a 21% relative improvement on EWER compared with the standard self-training approach.